Problem 1
Let f(x) = (a – 5)x^2 + (b – 3)x + c and g(x) = 3x^2 – 5x – 4.
If f(1) = g(1), f(-8) = g(-8), f(-79) = g(-79), find the value of a + 2b + 2c.
Problem 2
A video production company is planning to produce an instructional movie. The producer estimates that it will cost $250,000 to shoot the video and $150 per unit to copy and distribute the CD. The wholesale price of the disc is $750 per unit.
a. Write the cost function C, the revenue function R, and the profit function P.
b. How many CD units must be sold in order to have a profit? A loss? A break-even?

Answer by Theo(13342)

first proglem:

unless i'm missing something, i believe f(x) will be equal to g(x) when (a-5) = 3 and (b-3) = -2 and c = -4

when that happens, f(x) and g(x) will be identical.

a-5 = 3 when a = 8
b - 3 = -5 when b = -2
c = -4 when c = -4

you get a = 8, b = -2, c = -4

a + 2b + 2c is therefore equal to 8 + (-4) + (-8) which is equal to 8 - 12 which is equal to -4.

here's the graph of both equations.
this graph shows that the values of:

f(1) and g(1) are the same
f(-8) and g(-8) are the same.
f(-79) and g(-79) are the same.

first graph shows f(1) and f(-8) which are the same as g(1) and g(-8).

second graph shows f(-79) which is the same as g(-79).

$$$

$$$

second problem:

A video production company is planning to produce an instructional movie. The producer estimates that it will cost $250,000 to shoot the video and $150 per unit to copy and distribute the CD. The wholesale price of the disc is $750 per unit.
a. Write the cost function C, the revenue function R, and the profit function P.
b. How many CD units must be sold in order to have a profit? A loss? A break-even?

cost function is c = 250,000 + 150 * x

x = number of units.

revenue function is r = 750 * x

profit function is p = r - c

r is the revenue
p is the profit
x is the number of units produced and sold.
c is the cost.

break even is when the revenue equal the cost.

you will get r = c which then becomes:

750 * x = 250,000 + 150 * x

subtract 150 * x from both sides of this equation to get:

750 * x - 150 * x = 250,000

combine like terms to get:

600 * x = 250,000

solve for x to get:

x = 250,000 / 600 = 416 and 2/3 units.

he will break even when he produces and sells 416 and 2/3 units.

if he produces and sells less than that, he will lose money.

if he produces and sells more than that, he will make money.

examples:

he produces and sells 416 and 2/3 units.
revenue is 750 * 416 and 2/3 = 312,500.
cost is 250,000 + 150 * 416 and 2/3 = 312,500.
he broke even.

he produces and sells 200 units.
revenue is 750 * 200 = 150,000
cost is 250,000 + 150 * 200 = 280,000
he loses money.

he produces and sells 600 units.
revenue is 750 * 600 = 450,000
cost is 250,000 + 150 * 600 = 340,000
he makes money.

this is shown graphically below:

the blue line represents the cost.
the red line represents the revenue.

$$$

in formula terms:

break-even is when r = c
profit is when r > c
loss is when r < c

this leads to the following equations:

750*x = 250,000 + 150*x (break-even)
750*x > 250,000 + 150*x (make a profit)
750*x <= 250,000 + 150*x (take a loss)

i specifically said "produces and sells".

this is an important assumption.

it is possible that he can produce more than he sells.

that becomes a different type of problem that is more complex to solve.

i'm assuming you are not into those types of problems yet.

hence the assumption.